OpenStudy (anonymous):

For how many two-digit positive numbers will tripling the tens digits give us a two-digit number that is triple the original number

OpenStudy (anonymous):

Exactly three: 10, 20, and 30. Any number above 39 is to large, because 4*3=12 is greater than 9. Any number less than 10 is obviously too small. The decimal form is as follows: n=10a+b where a and b are whole numbers less than ten. If we are to have 10(3a)+b=3(10a+b), then b=0.

OpenStudy (anonymous):

Thank you